ABC Co. and XYZ Co. are identical firms in all respects except for their capital structure. ABC is all equity financed with $675,000 in stock. XYZ uses both stock and perpetual debt; its stock is worth $337,500 and the interest rate on its debt is 7.5 percent. Both firms expect EBIT to be $72,000. Ignore taxes.

  

a.

Rico owns $50,625 worth of XYZ’s stock. What rate of return is he expecting? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. Suppose Rico invests in ABC Co. and uses homemade leverage. Calculate his total cash flow and rate of return. (Do not round intermediate calculations and enter your rate of return answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What is the cost of equity for ABC and XYZ? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
d. What is the WACC for ABC and XYZ? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

A в Rico Net income Dividends received a) Rate of return с $ 46,688 |=72000-(675000-337500)*7.5% $ 7,003 =46688*(50625/337500

*Please rate thumbs up